The perimeter of a rectangle is 40cm and it's area is 84cm squared.
a). If the width of the rectangle is xcm, express the length of the rectangle in terms of x
b). Find the length and width of the rectangle

Answer by Cromlix(4381)

perimeter = 2*length + 2*width
Area = length * width
Length = y Width = x
Area = y * x
84 = y*x
Length (y) = 84/x
Substitute y = 84/x into
Perimeter = 2*y + 2*x
40 = 2*84/x + 2x
40 = 168/x + 2x
multiply through by x
40x = 168 + 2x^2
Rearrange
2x^2 - 40x + 168 = 0
Divide through by 2
x^2 - 20x + 84 = 0
(x - 6)(x - 14)
Therefore length = 6cm and width = 14 cm
OR length = 14cm and width = 6cm
Hope this helps
:-)